# O1-Pruner: Length-Harmonizing Fine-Tuning for O1-Like Reasoning Pruning

### v1

@@ Abstract
Recently, long-thought reasoning LLMs, such as OpenAI’s O1, adopt extended reasoning processes similar to how humans ponder over complex problems. This reasoning paradigm significantly enhances the model’s problem-solving abilities and has achieved promising results. However, long-thought reasoning process leads to a substantial increase in inference time. A pressing challenge is reducing the inference overhead of long-thought LLMs while ensuring accuracy. In this paper, we experimentally demonstrate that long-thought reasoning models struggle to effectively allocate token budgets based on problem difficulty and reasoning redundancies. To address this, we propose Length-Harmonizing Fine-Tuning (O1-Pruner), aiming at minimizing reasoning overhead while maintaining accuracy. This effective fine-tuning method first estimates the LLM’s baseline performance through pre-sampling and then uses RL-style fine-tuning to encourage the model to generate shorter reasoning processes under accuracy constraints. This allows the model to achieve efficient reasoning with lower redundancy while maintaining accuracy. Experiments on various mathematical reasoning benchmarks show that O1-Pruner not only significantly reduces inference overhead but also achieves higher accuracy, providing a novel and promising solution to this challenge. Our code is coming soon at https://github.com/StarDewXXX/O1-Pruner
@@ 1 Introduction
In this paper, we first revisit the long-thought reasoning processes. we observe that the reasoning processes in long-thought reasoning LLMs often exhibit significant redundancies, which leads to inefficient use of computational resources. This inefficiency not only increases inference costs but also highlights a fundamental limitation in the models’ ability to adapt their reasoning depth to suit the demands of diverse tasks. Building on this analysis, we formulate an optimization objective aimed at minimizing reasoning overhead while maintaining accuracy as a constraint. To achieve this objective, we propose a RL-based loss function. Our experiments are conducted using open-source long-thought reasoning LLMs, and we compare our approach against several competing methods like SFT and DPO (Rafailov et al., 2024). Through extensive experiments, we demonstrate the efficiency of our proposed methods. Additionally, we perform further studies on the influence of hyperparameters and dataset diffucilty on our approach, in order to gain deeper insights into the characteristics and behavior of this novel framework.
@@
  • We identify and experimentally validate the existence of length disharmony in the reasoning process of long thought models, which leads to redundant inference overhead.
@@ 2 Related Work
2.1 Chain of Thought
Chain-of-Thought (CoT) (Wei et al., 2023) prompting has emerged as a powerful technique for improving the reasoning capabilities of large language models (LLMs). By encouraging the model to solve complex problems step by step, CoT significantly enhances the accuracy and interpretability of its outputs. CoT is particularly effective for tasks that requiring multiple solving steps, such as mathematical problem-solving and logical reasoning. Beyond the basic CoT paradigm, many innovative frameworks like Tree of Thought (ToT) (Yao et al., 2023) and Graph of Thought (GoT) (Besta et al., 2024) expand upon the CoT architecture by investigating various reasoning trajectories or integrating structures based on networks.
Besides, Chain-of-thought reasoning also furnishes discernible traces of logical progression, enabling human to comprehend the model’s decision-making pathway, thereby rendering the reasoning process both transparent and credible.
2.2 Inference-time Scaling
@@
Inference-time scaling refers to the ability of large language models (LLMs) to improve their outputs by utilizing additional computation during inference time. Recent studies (Snell et al., 2024) have explored how scaling inference-time computation can enhance the performance of LLMs on challenging prompts. This approach draws parallels to human reasoning, where additional cognitive effort is often allocated to complex problems. In addition to increasing the number of candidate solutions or searching different steps, OpenAI’s O1 inference (OpenAI, 2024) demonstrates that extending the length of the solution generated during reasoning can also significantly enhance the model’s performance.
2.3 LLM Alignment
@@
LLM alignment (Shen et al., 2023) constitutes a technical process aimed at guaranteeing that the responses generated by large language models are not only precise and logically consistent but also secure, morally sound, and aligned with the expectations of both developers and users. Ensuring that these expansive language models are in harmony with human preference is crucial for leveraging their immense capabilities in a manner that is both reliable and conscientious. Common methodologies employed in LLM alignment include Supervised Fine-Tuning (SFT), Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022), and Direct Preference Optimization (DPO), among others. The discourse on long thought reasoning optimization presented in this paper can be regarded as an extended setting of LLM alignment, where human preferences are inclined towards shorter outputs (faster inference) and enhanced reasoning accuracy.
@@
Table 1: Accuracy-Length Relationship at Distribution Level. A larger interval number indicates a longer solution length. For both models, the average accuracy is higher when the solution length is short.
3 Length Disharmony in Long Thought Reasoning
We employ the term ”Length Disharmony” to characterize the phenomenon of inefficiency in the reasoning process of long-thought reasoning, when the model generates responses of varying lengths, among which the shorter responses possess sufficiently high accuracy, thereby rendering the longer responses a superfluous expenditure of computational resources. Besides, due to the quadratic complexity of the Transformer architecture, this will significantly leads to an increase in inference time.
From this, we can ascertain the relationship between accuracy and length at the instance level, which is shown in Fig 1. It is evident that there exists a markedly inconsistent relationship between length and accuracy across different problems. The highest accuracy may manifest within the shortest, intermediate, or longest length intervals. Particularly, we observe that a relatively high accuracy is maintained even within the intervals of shorter lengths.
Furthermore, by calculating the average accuracy across all problems within different intervals, we have derived the relationship between accuracy and length at the distribution level, which is shown in Table 1. At the distribution level, we observe that the average accuracy is still high when the response length is short.
Therefore, we can conclude that long-thought models exhibit a characteristic of length disharmony during reasoning, which leads to redundant computational overhead in the inference phase. This reasoning redundancy can be mitigated, as high accuracy is still maintained even at shorter lengths. From this perspective, we propose Length-Harmonizing Fine-Tuning to optimize long-thought reasoning, enabling it to maintain high accuracy while reducing inference redundancy.
@@ 4 Methodology
In this section, we elaborate on our proposed Length-Harmonizing Fine-Tuning in detail and provide a simple and intuitive mathematical analysis elucidating how our method works for optimize long thought of reasoning.
@@ 4.1 Problem Setup
We consider a Large Language Model (LLM) parameterized by $\mathbf{\theta}$ and denoted as $\pi_{\mathbf{\theta}}$. In the context of math problem solving, the LLM accepts a sequence $\mathbf{x}=[x_{1},\ldots,x_{n}]$, commonly termed as the problem, and then generate a corresponding solution $\mathbf{y}=[y_{1},\ldots,y_{m}]$. Hence, the solution $\mathbf{y}$ is construed as a sample drawn from the conditional probability distribution $\pi_{\mathbf{\theta}}(\cdot|\mathbf{x})$. The conditional probability distribution $\pi_{\mathbf{\theta}}(\mathbf{y}|\mathbf{x})$ can be decomposed as follows:
$\pi_{\mathbf{\theta}}(\mathbf{y}|\mathbf{x})=\prod_{j=1}^{m}\pi_{\mathbf{\theta}}(y_{j}|\mathbf{x},\mathbf{y}_{<j}),$ (1)
@@
$\max_{\pi_{\theta}}\mathbb{E}_{(x,y)\sim\mathcal{D}}\Big{[}\log\pi_{\mathbf{\theta}}(\mathbf{y}\ |\ \mathbf{x})\Big{]}$ (2)
4.2 Length-Harmonizing Fine-Tuning
To start with, let’s assume that $\pi_{\theta}$ is a LLM that can solve math problems with long thought with redundancy and disharmony. we hypothesize that the reasoning paths represented by output thought of language model $\pi_{\theta}$ contain redundancies and lack proper coordination. To address this, we propose an optimization objective that ensures no degradation in accuracy while tackling the issue from two perspectives. First, at the overall level, we aim to shorten the reasoning paths. Second, we encourage the model to output shorter answers for simpler problems, while for more complex problems, we guide the model to learn the correct reasoning paths, which, according to the inference scaling law, typically involve longer reasoning sequences. Consider a problem $x$. We define $L(y)$ to be the length (counted by token) of solution $y$. Considering a reference model $\pi_{ref}$, we hope to to reduce the solution length of the policy model relative to that of the reference model, which can be formulated as:
@@ 4.2 Length-Harmonizing Fine-Tuning
$\mathbb{E}_{x\sim D,y\sim\pi_{\theta}(y|x)}A(x,y)\geq\mathbb{E}_{x\sim D,y^{\prime}\sim\pi_{ref}(y^{\prime}|x)}A(x,y^{\prime})$ (4)
Therefore, we can establish our optimization objective as:
$\max\mathbb{E}_{x\sim D}\left[\mathbb{E}_{y\sim\pi_{\theta}(y|x),y^{\prime}\sim\pi_{ref}(y|x)}\frac{L(y^{\prime})}{L(y)}-1\right]$
s.t. $\mathbb{E}_{x\sim D,y\sim\pi_{\theta}(y|x)}A(x,y)\geq\mathbb{E}_{x\sim D,y^{\prime}\sim\pi_{ref}(y^{\prime}|x)}A(x,y^{\prime})$ (5)
To solve this constrained optimization problem, we employ the Lagrange multipliers, incorporating constraint into the objective function as a penalty term. Specifically, the constraint on accuracy is added to the objective with a penalty weight $\lambda\geq 0$:
@@
$\lambda(A(x,y)-\mathbb{E}_{y^{\prime}\sim\pi_{ref}(y^{\prime}|x)}A(x,y^{\prime}))$ (7)
@@
$\bar{L}_{ref}(x)=\sum_{y^{\prime}_{i}\sim\pi_{ref}(\cdot|x)}^{K}L(y^{\prime}_{i})$ (8)
$\bar{A}_{ref}(x)=\sum_{y^{\prime}_{i}\sim\pi_{ref}(\cdot|x)}^{K}A(x,y^{\prime}_{i})$ (9)
So that our objective can be re-written as:
@@
$L^{\text{LH}}(\theta)=-\mathbb{E}_{x\sim D,y\sim\pi_{ref}(y|x)}[\min(r(\theta)R_{LH}(x,y),\,$
$\text{clip}(r(\theta),1-\epsilon,1+\epsilon)R_{LH}(x,y))]$ (11)
, where $r(\theta)=\frac{\pi_{\mathbf{\theta}}(\mathbf{y}|\mathbf{x})}{\pi_{ref}(\mathbf{y}|\mathbf{x})}$. clip() is the clipping function of PPO.
@@ 5.1 Experiment Setup
5.1.1 Model
The long thought models we chosen for our experiment are Marco-o1-7B and QwQ-32B-Preview, which demonstrate excellent performance on a wide range of math problem-solving tasks. For Marco-o1-7B, we utilize full-parameter fine-tuning; however, for the larger-scale QwQ-32B-Preview, our computational resources are not able to support full-parameter training. As a result, we adopt Parameter-Efficient Fine-Tuning (Han et al., 2024). After evaluating both LoRA (Hu et al., 2021) and Freeze Fine-Tune, we observed that Freeze Fine-Tune yields much better performance. Therefore, we selected this fine-tuning approach for our experiments.
5.1.2 Dataset
The dataset used for training is MATH. It comprises approximately 10k math problem of high school level accompanied with both ground truth solution and ground truth answer. Since the ground truth solution is not need for our experiment, we only use the problem-answer pairs. For training, we selected 5,000 problems from the MATH Trainset. For Marco-o1-7B, we generated 16 solutions for each problem; for QwQ-32B-Preview, we generated 12 solutions for each problem.
The dataset utilized for testing encompasses the test sets of MATH, GSM8k (Cobbe et al., 2021), and GaoKao (mathematical) (Zhang et al., 2024), comprising a diverse range of mathematical problems with varying levels of difficulty.
5.1.3 Competing Methods
To validate the superiority of our method for long thought reasoning optimization tasks, we have selected the following comparative methods:
  • Fast-Solving Prompt The Fast-Solving Prompt is a prompting technique wherein we instruct the model within the prompt to solve the given problem as swiftly as possible, aiming to achieve the desired reduction in reasoning length.
  • SFT For the Supervised Fine-Tuning (SFT) method, we curated the training dataset by selecting the two shortest correct solutions for each problem, ensuring that the model is exposed to examples that embody both accuracy and conciseness. These solutions were then used to train the model following the standard SFT pipeline.
  • DPO For the implementation DPO, we meticulously selected two of the shortest correct solutions to serve as the chosen samples, which exemplify efficiency and precision in problem-solving. Conversely, to represent the reject sample, we opted for the longest solution available.
5.1.4 Evaluation Metric
In our experiments, we employed average accuracy, average length and Accuracy-Efficiency Score (AES) as key metrics.
@@ 5.1.4 Evaluation Metric
By analyzing these three metrics, we can assess whether the model achieves a desirable balance between reasoning accuracy and reasoning length.
5.2 Results
The experimental results, summarized in Table 2, demonstrate the performance of various methods across different evaluation metrics. The proposed O1-Pruner consistently achieves superior performance in balancing reasoning accuracy and efficiency compared to baseline and competing methods. Notably, it exhibits the best trade-off between accuracy and reasoning length across all datasets, as further supported by its significantly higher Accuracy-Efficiency Score (AES) values.

### v2

@@ Abstract
Recently, long-thought reasoning LLMs, such as OpenAI’s O1, adopt extended reasoning processes similar to how humans ponder over complex problems. This reasoning paradigm significantly enhances the model’s problem-solving abilities and achieves promising results. However, long-thought reasoning process leads to a substantial increase in inference time. A pressing challenge is reducing the inference overhead of long-thought LLMs while ensuring accuracy. In this paper, we identify that long-thought reasoning models struggle to effectively allocate token budgets based on problem difficulty and reasoning redundancies. To address this, we propose Length-Harmonizing Fine-Tuning (O1-Pruner), aiming at minimizing reasoning overhead while maintaining accuracy. This effective fine-tuning method first estimates the LLM’s baseline performance through pre-sampling and then uses RL-style fine-tuning to encourage the model to generate shorter reasoning processes under accuracy constraints. This allows the model to achieve efficient reasoning with lower redundancy while maintaining accuracy. Experiments on various mathematical reasoning benchmarks show that O1-Pruner not only significantly reduces inference overhead but also achieves higher accuracy, providing a novel and promising solution to this challenge. Our code is coming soon at https://github.com/StarDewXXX/O1-Pruner
@@ 1 Introduction
In this paper, we first revisit the long-thought reasoning processes. we observe that the reasoning processes in long-thought reasoning LLMs often exhibit significant redundancies, which leads to inefficient use of computational resources. This inefficiency not only increases inference costs but also highlights a fundamental limitation in the models’ ability to adapt their reasoning depth to suit the demands of diverse tasks. Building on this analysis, we formulate an optimization objective aimed at minimizing reasoning overhead while maintaining accuracy as a constraint. Our approach introduces a Length-Harmonizing Reward, which explicitly rewards shorter solutions while penalizing accuracy degradation. By embedding this reward into a RL-based framework, we enable the model to optimize for efficiency without compromising performance. Moreover, our method incorporates an off-policy training strategy inspired by Proximal Policy Optimization (PPO), which aimed at reducing training complexity while maintaining robustness.
Our experiments are conducted using open-source long-thought reasoning LLMs, and we compare our approach against several competing methods like SFT and DPO (Rafailov et al., 2024). Through extensive experiments, we demonstrate the efficiency of our proposed methods. Additionally, we perform further studies on the influence of hyperparameters and dataset difficulty on our approach, in order to gain deeper insights into the characteristics and behavior of this novel framework.
@@
  • We design a simple experiment and identify a critical issue in the reasoning process of long-thought models, referred to as length disharmony, which leads to redundant inference overhead.
@@ 2 Related Work
Inference-time Scaling. Inference-time scaling refers to the ability of large language models (LLMs) to improve their outputs by utilizing additional computation during inference time. Recent studies (Snell et al., 2024) have explored how scaling inference-time computation can enhance the performance of LLMs on challenging prompts. This approach draws parallels to human reasoning, where additional cognitive effort is often allocated to complex problems. In addition to increasing the number of candidate solutions or searching different steps, OpenAI’s O1 inference (OpenAI, 2024) demonstrates that extending the length of the solution generated during reasoning can also significantly enhance the model’s performance.
@@
LLM Alignment. LLM alignment (Shen et al., 2023) constitutes a technical process aimed at guaranteeing that the responses generated by large language models are not only precise and logically consistent but also secure, morally sound, and aligned with the expectations of both developers and users. Ensuring that these expansive language models are in harmony with human preference is crucial for leveraging their immense capabilities in a manner that is both reliable and conscientious. Common methodologies employed in LLM alignment include Supervised Fine-Tuning (Zhou et al., 2023), Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022), and Direct Preference Optimization (DPO), among others. The discourse on long thought reasoning optimization presented in this paper can be regarded as an extended setting of LLM alignment, where human preferences are inclined towards shorter outputs (faster inference) and enhanced reasoning accuracy.
CoT Compression. Chain-of-Thought (CoT) (Wei et al., 2023) and its variations (ToT, (Yao et al., 2023), GoT (Besta et al., 2024)) are powerful techniques for improving the reasoning capabilities of LLMs. Although CoT is highly effective, it introduces additional computational overhead. Consequently, several studies have attempted to address this issue. For example, (Han et al., 2024a) introduced a token-budget-aware reasoning framework for large language models (LLMs), which dynamically allocates token budgets according to the complexity of different problems and leverages these budgets to guide the reasoning process. C3oT (Kang et al., 2024) employs GPT-4 as a compressor to retain critical information during the reasoning process, thereby reducing reasoning redundancy. Furthermore, several approaches try to utilize continuous representations to mitigate the computational overhead associated with Chain-of-Thought (CoT). For example, CCoT (Cheng & Durme, 2024) reduces reasoning overhead by generating contentful and continuous contemplation tokens of variable sequence lengths. COCONUT (Hao et al., 2024) train LLMs to reason with fewer thinking tokens during inference in a continuous latent space. However, unlike traditional approaches that focus on compressing normal Chain-of-Thought (CoT), our method centers on long thought reasoning and reduces redundancy in such reasoning by optimizing the reasoning paths instead of compressing each reasoning step.
Some concurrent works, such as (Chen et al., 2024), have identified the issue of overthinking in O1 reasoning and employs SimPO(Meng et al., 2024) for optimization, which is based on the view of preference learning. And (Team et al., 2025) propose long2short RL, using long-CoT techniques to improve short-CoT models. However, in this paper we analyze the long-thought model from a different perspective of length distribution. Moreover, we establish an optimization problem and propose a RL-based method to optimize the model, which provides a different and novel perspective for subsequent research.
@@
3 Revisiting the “Length Disharmony” in Long Thought Reasoning
We employ the term “Length Disharmony” to characterize the phenomenon of inefficiency in the reasoning process of long-thought reasoning, when the model generates responses of varying lengths, among which the shorter responses possess sufficiently high accuracy, thereby rendering the longer responses a superfluous expenditure of computational resources. Besides, due to the quadratic complexity of the Transformer architecture, this will significantly leads to an increase in inference time.
Accuracy-Length Relationship at Instance Level. From the data we collected, we can ascertain the relationship between accuracy and length at the instance level, which is shown in Figure 1. It is evident that there exists a markedly inconsistent relationship between length and accuracy across different problems. The highest accuracy may manifest within the shortest, intermediate, or longest length intervals. Specifically, we observe that relatively high accuracy is preserved even within shorter-length intervals.
Accuracy-Length Relationship at Distribution Level. Furthermore, by calculating the average accuracy across all problems within different intervals, we have derived the relationship between accuracy and length at the distribution level, which is shown in Table 1. At the distribution level, our analysis reveals a consistent trend where shorter response lengths are associated with higher average accuracy rates. This observation can be explained by the premise that a shorter response length typically signifies the model’s ability to identify the optimal solution path more efficiently, consequently requiring fewer iterative processes of reflection and backtracking.
Therefore, we can conclude that long-thought models exhibit a phenomenon of length disharmony during reasoning, which leads to redundant computational overhead in the inference phase. This reasoning redundancy can be mitigated, as high accuracy is still maintained even at shorter lengths. From this perspective, we propose Length-Harmonizing Fine-Tuning (O1-Pruner) to optimize long-thought reasoning, enabling it to maintain high accuracy while reducing inference redundancy.
Table 1: Accuracy-Length Relationship at Distribution Level. A larger interval number indicates a longer solution length. The average accuracy is higher when the solution length is short.
@@ 4 Methodology
In this section, we elaborate on our proposed Length-Harmonizing Fine-Tuning (O1-Pruner) in detail and provide a simple and intuitive mathematical analysis elucidating how our method works for optimize long thought of reasoning.
@@ 4.1 Problem Setup
We consider a LLM parameterized by $\mathbf{\theta}$ and denoted as $\pi_{\mathbf{\theta}}$. In the context of math problem solving, the LLM accepts a sequence $\mathbf{x}=[x_{1},\ldots,x_{n}]$, commonly termed as the problem, and then generate a corresponding solution $\mathbf{y}=[y_{1},\ldots,y_{m}]$. Hence, the solution $\mathbf{y}$ is construed as a sample drawn from the conditional probability distribution $\pi_{\mathbf{\theta}}(\cdot|\mathbf{x})$. The conditional probability distribution $\pi_{\mathbf{\theta}}(\mathbf{y}|\mathbf{x})$ can be decomposed as follows:
$\pi_{\mathbf{\theta}}(\mathbf{y}|\mathbf{x})=\prod_{j=1}^{m}\pi_{\mathbf{\theta}}(y_{j}|\mathbf{x},\mathbf{y}_{<j}).$ (1)
@@
$\max_{\pi_{\theta}}\mathbb{E}_{(x,y)\sim\mathcal{D}}\Big{[}\log\pi_{\mathbf{\theta}}(\mathbf{y}\ |\ \mathbf{x})\Big{]}.$ (2)
4.2 Length-Harmonizing Fine-Tuning (O1-Pruner)
To start with, let’s assume that $\pi_{\theta}$ is a LLM that can solve math problems with long thought with redundancy and disharmony. we hypothesize that the reasoning paths represented by output thought of language model $\pi_{\theta}$ contain redundancies and lack proper coordination. To address this, we propose an optimization objective that ensures no degradation in accuracy while tackling the issue from two perspectives. First, at the overall level, we aim to shorten the reasoning paths. Second, we encourage the model to output shorter answers for simpler problems, while for more complex problems, we guide the model to learn the correct reasoning paths, which, according to the inference scaling law, typically involve longer reasoning sequences. Given a problem $x$, we define $L(y)$ as the length (counted by token) of the solution $y$. Considering a reference model $\pi_{ref}$, we reduce the solution length of the policy model relative to that of the reference model, which can be formulated as:
$\max\mathbb{E}_{x\sim D}\left[\mathbb{E}_{y\sim\pi_{\theta}(y|x),y^{\prime}\sim\pi_{ref}(y|x)}\frac{L(y^{\prime})}{L(y)}-1\right].$ (3)
$\!\!\!\mathbb{E}_{x\sim D,y\sim\pi_{\theta}(y|x)}A(x,y)\geq\mathbb{E}_{x\sim D,y^{\prime}\sim\pi_{ref}(y^{\prime}|x)}A(x,y^{\prime}).$ (4)
Therefore, we can establish our optimization objective as:
@@ 4.2 Length-Harmonizing Fine-Tuning (O1-Pruner)
s.t. $\mathbb{E}_{x\sim D,y\sim\pi_{\theta}(y|x)}A(x,y)\geq\mathbb{E}_{x\sim D,y^{\prime}\sim\pi_{ref}(y^{\prime}|x)}A(x,y^{\prime}).$
To solve this constrained optimization problem, we incorporate constraint into the objective function as a penalty term. Specifically, the constraint on accuracy is added to the objective with a penalty weight $\lambda\geq 0$:
@@
$\lambda(A(x,y)-\mathbb{E}_{y^{\prime}\sim\pi_{ref}(y^{\prime}|x)}A(x,y^{\prime})).$ (7)
@@
$\bar{L}_{ref}(x)=\frac{1}{K}\sum_{i=1}^{K}L(y^{\prime}_{i}),\quad y^{\prime}_{i}\sim\pi_{ref}(\cdot\mid x);$ (8)
$\bar{A}_{ref}(x)=\frac{1}{K}\sum_{i=1}^{K}A(x,y^{\prime}_{i}),\quad y^{\prime}_{i}\sim\pi_{ref}(\cdot\mid x);$ (9)
This approach is widely employed in Policy Gradient with Baseline. Furthermore, a recently proposed method GRPO (Shao et al., 2024) adopts a similar technique to reduce training overhead. Based on this technique, our objective can be approximated as:
@@
$L^{\text{LH}}(\theta;x,y)=-\mathbb{E}_{x\sim D,y\sim\pi_{ref}(y|x)}\big{[}\min(r(\theta)R_{LH}(x,y),\,$
$\qquad\qquad\text{clip}(r(\theta),1-\epsilon,1+\epsilon)R_{LH}(x,y))\big{]},$ (11)
where $r(\theta)=\frac{\pi_{\mathbf{\theta}}(\mathbf{y}|\mathbf{x})}{\pi_{ref}(\mathbf{y}|\mathbf{x})}$. clip() is the clipping function.
@@ 4.3 Understanding the Loss Function
To the end, we summarize the training procedure of our proposed O1-Pruner in Algorithm 1.
Algorithm 1 O1-Pruner
1:  Input: LLM $\pi_{\theta}$, Dataset $\mathcal{D}=\{(x^{i},a^{i})\}_{i\in[N]}$
2:  Initialize: $\pi_{ref}=\pi_{\theta}$
3:  for $i=1$ to $N$ do
4:     sampling K solutions $y^{i}_{1}$, …,$y^{i}_{K}$ from $\pi_{ref}(\cdot|x_{i})$
5:     calculating $\bar{L}_{ref}(x^{i})=\frac{1}{K}\sum_{k=1}^{K}L(y^{i}_{k})$
6:     calculating $\bar{A}_{ref}(x^{i})=\frac{1}{K}\sum_{k=1}^{K}A(x^{i},y^{i}_{k})$
7:     randomly select m ($m\leq K$) solutions from $y^{i}_{1}$, …,$y^{i}_{K}$
8:     Update $\theta=\mathop{\arg\min}\limits_{\theta}\sum_{j=1}^{m}L^{\text{LH}}(\theta;x^{i},y^{i}_{j})$
9:  end for
10:  Output: Updated LLM $\pi_{\theta}$
@@ 5 Experiments
In this section, we conduct extensive experiments to verify the efficacy of the proposed O1-Pruner.
@@ 5.1 Experiment Setup
Long-thought Models. The long thought models we chosen for our experiment are Marco-o1-7B and QwQ-32B-Preview, which have demonstrated excellent performance on a wide range of math problem-solving tasks. For Marco-o1-7B, we utilize full-parameter fine-tuning; however, for the larger-scale QwQ-32B-Preview, our computational resources are not able to support full-parameter training. As a result, we adopt Parameter-Efficient Fine-Tuning (Han et al., 2024b). After evaluating both LoRA (Hu et al., 2021) and Freeze Fine-Tune, we observed that Freeze Fine-Tune yields much better performance. Therefore, we selected this fine-tuning approach for our experiments.
Dataset. The dataset used for training is MATH. It comprises approximately 10k math problem of high school level accompanied with both ground truth solution and ground truth answer. Since the ground truth solution is not need for our experiment, we only use the problem-answer pairs. For training, we selected 5,000 problems from the MATH Trainset. For Marco-o1-7B, we generated 16 solutions for each problem; for QwQ-32B-Preview, we generated 12 solutions for each problem. The dataset utilized for testing encompasses the test sets of MATH, GSM8k (Cobbe et al., 2021), and GaoKao (mathematical) (Zhang et al., 2024), comprising a diverse range of mathematical problems with varying levels of difficulty.
Baselines. To validate the superiority of our method for long thought reasoning optimization tasks, we have selected the three representatively comparative methods. (i) Fast-Solving Prompt: The Fast-Solving Prompt is a prompting technique wherein we instruct the model within the prompt to solve the given problem as swiftly as possible, aiming to achieve the desired reduction in reasoning length. (ii) SFT: For the SFT method, we curated the training dataset by selecting the two shortest correct solutions for each problem, ensuring that the model is exposed to examples that embody both accuracy and conciseness. These solutions were then used to train the model following the standard SFT pipeline. (iii) DPO: For the implementation of DPO, we meticulously selected two of the shortest correct solutions to serve as the chosen samples, which exemplify efficiency and precision in problem-solving. Conversely, to represent the reject sample, we opted for the longest solution available.
Evaluation Metric. We employ the following average accuracy, average length and Accuracy-Efficiency Score (AES) as key metrics to assess whether the model achieves a desirable balance between reasoning accuracy and length:
@@
5.2 Experimental Results